Prev Next

SW&IT Meeting Minutes 2020-09-14

Attendees: Carl, David, Igor, Matthieu, Stephen, Trey, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200907
LR20200907 - Bec, Matthieu 32.0
LR20200907 - Holmberg, Carl 40.0
LR20200907 - Edmeades, David 32.0
LR20200907 - Sola, Igor 40.0
LR20200907 - Hooper, Stephen 0.0
LR20200907 - Bryant, Trey 0.0
LR20200907 - Zhang, Yang 16.0
LR20200907 - Zhang, Xianyu 32.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

IS

Last week:

  • PIT
    • Most of the week dedicated to this work. I had to find new logic in order to include PEPSI as a pairing instrument w/ binocular observations. Done a special release to test new logic

This week:

  • Safety Training
  • PIT
    • Work on the backend admin tool
  • LUCI
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
    • Finish document and configuration for WebIO

DE

  • VPN accounts
    • Sort out procedure re: AUP
    • Do we care about how many we create?
  • JPH printer order
  • Receipt of second RMA HD
  • ATT Zoom accounts
    • No complaints so far, though unsure of how much they've been tested so far
  • Sorted out VPN Affiliates group DNS issue with SH
  • Campus Covid training.

Next:
  • Steward visit to swap HD/bad sector reallocations
    • Any other IT business while I'm in?
  • Work on Vault backup config wrt storage drive, dealing with VMWare not cleaning up VM disks
  • Cisco training

CH

  • Last week:
    • indi-core: updated pyindi.
    • ahs: added LBTI alert client.
  • This week:

YZ

  • Previous:
    • Some AO-Cravat/OSCO and TMS work.
  • Next:
    • Upcoming TMS test (tentative on Wednesday).
    • Usual OSCO/Cravat and TMS work.
    • Some LBC CMU clean-up.
    • Several issue tracks popped up (IT#8176: RB_Science window problem, IT#4566: LBC red rotator communication error, IT#8182: PMC GUI crash, IT#8179: PCS co-pointing limit problems, IT#3949: PCS rotator time-to-limit predication error).
    • Archive file checksum follow-up?
    • Fixing tcs-test for Alex Becker.
    • UAccess COVID-19 safety training.

MB

  • Last week
    • OC, Management
    • ATT KoM, references document, 1st drast plan
    • DIMM investigations
    • Review SHARK-VIS RFAs
    • meetings Archive, etc.
    • AAS job regiser ad
  • This week
    • ATT TOC, IssueTrack review, 1st draft
    • ROWG restart

SH

  • DIMM hardware
    • old hardware, new machine
    • replacement 2950
    • PCI card and new HDD
  • DIMM software
    • CentOS 6.10 compilation - Ice log issues
    • Driver issues
    • Find missing files for EDT driver, try different EDT drivers
  • VPN
    • DNS issues resolved for Affiliates
    • Add a couple of external sites to split-tunneling
next:

  • DIMM Kernel panic on EDT card driver configuration
  • DIMM software issues, work with Leroy and hardware

XZ

  • Last week
    • Setup script for moving the AdSec and WFS log/telemetry files
    • "Clean" the AO log/telemetry files
    • SX AdSec small "dump" issue
    • Safety training
    • AO support
    • Meetings
  • This week
    • Handle shell rested but dome open in AOS?
    • Safety training
    • AO support

TB

  • Last week
    • Fixing the RMS functionality; pick chunk with highest mean value instead of chunk with highest RMS (misunderstood at first)
    • Redo report with figures now that the above ^ issue is resolved
  • This week
    • Work more on report?
    • Other next steps
Topic revision: r10 - 14 Sep 2020, YangZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback